# Enhancing the Performance of Aspect-Based Sentiment Analysis Systems

### v2

@@ 5 Conclusion
This study enhanced the performance of aspect-based sentiment analysis systems, called SentiSys, by leveraging an edge-enhanced bidirectional GCN. The SentiSys model employed Bi-LSTM to acquire a deep understanding of the syntax and semantics within the text. The hidden layer features derived from Bi-LSTM are utilized for constructing syntactic dependency trees in the Bi-GCN. Furthermore, SentiSys incorporated a transformer encoder layer designed to extract global information from the text, enhancing SentiSys’s ability to capture nuanced features in lengthy textual data. The dependency parsing layer within Bi-GCN extracts the graph structure based on the text’s syntax, facilitating the learning of connections between aspects and context. To further refine the model, an aspect-specific masking layer is implemented, reducing redundant information in the GCN’s hidden layer and ultimately improving the accuracy of the aspect-based sentiment analysis system. Finally, SentiSys combined the output from Bi-GCN and transformer, both of which serve as inputs for the sentiment classification classifier.
The SentiSys model faces two primary limitations: Firstly, the computation of the SDI, a noteworthy contribution of this study, was performed statistically. In future work, we aim to implement trainable vectors within the network to enrich the features associated with each edge. Secondly, the use of the transformer for feature enhancement prompts a discussion about the most effective stage for merging global information into the model. In future work, we aim to overcome the first limitation by enhancing the SentiSys model through the integration of trainable vectors within the network. This improvement is anticipated to facilitate a more dynamic and context-specific computation of the SDI, thereby enriching the features associated with each edge. Addressing the second limitation, we will delve into optimal strategies for effectively integrating global information into the model. We plan to explore techniques that maximize the transformer’s impact at specific stages, ensuring its contributions are strategically incorporated to enhance overall classification accuracy.

### v3

@@ 5 Conclusion
This study improved aspect-based sentiment analysis with SentiSys, using an edge-enhanced bidirectional GCN. Bi-LSTM provided deep text understanding, with hidden features used to construct syntactic dependency trees in Bi-GCN. SentiSys included a transformer encoder layer for global information extraction, enhancing feature capturing in lengthy text. The dependency parsing layer within Bi-GCN extracted graph structure based on text syntax, aiding aspect-context connections learning. An aspect-specific masking layer refined the model, reducing redundant information in GCN’s hidden layer and enhancing accuracy. Finally, SentiSys combined Bi-GCN and transformer outputs for sentiment classification. Two primary limitations were identified: statistical computation of SDI and transformer’s feature enhancement stage. Future work aims to address these limitations by integrating trainable vectors for dynamic SDI computation and exploring optimal strategies for transformer integration to enhance classification accuracy.
Acknowledgements
This research was supported by the National Natural Science Foundation of China (Grant No. 62402022), the International Research Fellow of Japan Society for the Promotion of Science (Postdoctoral Fellowships for Research in Japan [Standard]), and KAKENHI (20K11830) Japan.